OpenStudy (anonymous):

A spherical balloon is being inflated by a machine that increases the balloon's volume at a steady 1000cc per second. Please show that the balloon's radius after 6.0 sec of inflation is approximately 11.273cm

OpenStudy (superhelp101):

u have to use the formula dV/dt = 100pi

OpenStudy (superhelp101):

volume of sphere is V= 4/3 pi r^3 dV/dr = 4 pi r^2

OpenStudy (superhelp101):

dr/dt = 100pi / 4pi r^2 = 25/ r^2 then you ought to substitute r= 5ft dr/dt = 25/ 25= 1 ft/min

OpenStudy (superhelp101):

the radius is 5ft, Surface area, A =4pi r^2 dA/dr = 8pi r soo... To get dA/dt = dA/dr * dr/dt dA/dt = 8 pi r * 1 = 8 pi r then you have to substitute r = 5 ft. dA/dt = 40pi ft^2 /min

OpenStudy (superhelp101):

@roger97c u got it ?